OpenStudy (anonymous):

A mass of 4.00 kilograms is completely converted into energy. What is this amount of energy in joules? (1 J = 1 kg m2/s2) (Points : 3) 1.20 x 10^9 J 1.33 x 10^-8 J 3.60 x 10^17 J 4.4 x 10^-16 J

OpenStudy (anonymous):

Is it the last one?

OpenStudy (anonymous):

can you show your work please i know you have to use E=mc^2 i just am having trouble

OpenStudy (anonymous):

@RANE

OpenStudy (rane):

ok whts c ?

OpenStudy (rane):

use your formula sheet

OpenStudy (anonymous):

Thats what i need help with and my teacher never gave me one

OpenStudy (anonymous):

Im having a hard time with my teacher no one understands her

OpenStudy (rane):

ok w8

OpenStudy (rane):

ok c = 3.7 × 10–14 J and that is the speed of light

OpenStudy (rane):

so since u have both m and c u can find e

OpenStudy (anonymous):

where are you getting 3.7 x 10-^14?

OpenStudy (rane):

u can find it from internet, thats the speed of light, u can find it anywhere or ask anyone

OpenStudy (anonymous):

Ohh okay

OpenStudy (anonymous):

so now whats the next step

OpenStudy (anonymous):

Find E?

OpenStudy (rane):

yes E = MC^2 = (4)(3.7 x 10-^14)^2 = now just calculate

OpenStudy (anonymous):

so 4 * 3.7 is 14.8

OpenStudy (anonymous):

and then... how do you muliply 4 by 10^-14?

OpenStudy (rane):

i don't have the calculator with me right now wht did u get on your calculator ?

OpenStudy (anonymous):

Its not working in my calculator it keeps saying error :/ ughhh

OpenStudy (rane):

ok w8

OpenStudy (anonymous):

it says 5.476 E 29 when i dont put the (-) on the exponent (14)

OpenStudy (rane):

ok i did and i got 4.4 x 10^-16 J

OpenStudy (anonymous):

Okay just so i know how did you do it? lol

OpenStudy (anonymous):

Like calculate this (4)(3.7 x 10-^14)^2

OpenStudy (anonymous):

?

OpenStudy (rane):

yes

OpenStudy (rane):

hold on the internet is telling me the wrong sped of lich © the actual c =3*10^8 and then u get C as an answer

OpenStudy (anonymous):

Ugh lol im soooo confused haha how?

OpenStudy (rane):

ok just use value of c c =3*10^8

OpenStudy (anonymous):

Hmmm.. soo E = m(3*10^8)^2

OpenStudy (anonymous):

@RANE :(

OpenStudy (rane):

yessssss

OpenStudy (anonymous):

@Redenbacher1 did u get ur answer or need a help??? :)

OpenStudy (rane):

just insert the value of c and the mass in the formula E = mc^2

OpenStudy (anonymous):

@sarah786 and @RANE so its 3.60 x 10^17 J??

OpenStudy (rane):

yes thats is the answer

OpenStudy (anonymous):

yh u r right :)